OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined s^3-2s/s^2-64

OpenStudy (hoblos):

expression is undefined for s^2-64=0 (s-8)(s+8)=0 s-8=0 OR s+8=0 s=8 s=-8